# Legally Binding but Unfair? Towards Assessing Fairness of Privacy Policies

### v1

@@
Abstract
Privacy policies are expected to inform data subjects about their data protection rights. They should explain the data controller’s data management practices, and make facts such as retention periods or data transfers to third parties transparent. Privacy policies only fulfill their purpose, if they are correctly perceived, interpreted, understood, and trusted by the data subject. Amongst others, this requires that a privacy policy is written in a fair way, e.g., it does not use polarizing terms, does not require a certain education, or does not assume a particular social background.
In this work-in-progress paper, we outline our approach to assessing fairness in privacy policies. To this end, we identify from fundamental legal sources and fairness research, how the dimensions informational fairness, representational fairness and ethics / morality are related to privacy policies. We propose options to automatically assess policies in these fairness dimensions, based on text statistics, linguistic methods and artificial intelligence. Finally, we conduct initial experiments with German privacy policies to provide evidence that our approach is applicable. Our experiments indicate that there are indeed issues in all three dimensions of fairness. For example, our approach finds out if a policy discriminates against individuals with impaired reading skills or certain demographics, and identifies questionable ethics. This is important, as future privacy policies may be used in a corpus for legal artificial intelligence models.
Index Terms:
@@ Abstract
I Introduction
The General Data Protection Regulation (GDPR) [1] requires any organization, that manages personal data, to publish a privacy policy. A policy should make transparent how personally identifiable information is collected, shared, or used [2]. Privacy policies fulfill an important social task: They balance the information deficit between the holder and the subject of personal data, and create trust [3].
To ensure that privacy policies are properly understood, perceived, and accepted, they must be written in a fair way. This includes multiple dimensions. Informational Fairness is not just about what information is communicated, but also how it is communicated [4, 5, 6, 7, 8]. A negative example would be a policy consisting of complex legal phrases, which discriminates against people with dyslexia and non-native speakers. Representational Fairness is about biases towards certain groups represented in texts [9, 10, 11]. For example, a policy could discriminate against women by only using male word forms. Ethics and Morality are at the core of fairness [12, 13] and to some extent captured by legal frameworks [14, 1]. A negative example would be a policy, that tries to take exclusive rights to utilize the user’s personal data. However, many different definitions of fairness exist [15, 16], some of them are contradictory [17, 18], and to the best of our knowledge, none of them is tailored for privacy policies. It is also unclear, how well methods from natural language processing or artificial intelligence allow evaluating privacy policies for such issues.
Thus, our research question is as follows: How can we automatically assess informational fairness, representational fairness, and ethics / morality of privacy policies?
In this work-in-progress paper, we relate prominent definitions and concepts of fairness and bias to privacy policies. We propose approaches for automatically evaluating privacy policies, and we test this with selected policies. We make three contributions:
@@ I Introduction
Paper structure: Section II reviews related work. In Section III, we derive our problem statement from legal requirements. The Sections IV, V and VI explain how three dimensions of fairness can be automatically assessed. Section VII summarizes potential applications and use cases of our approach. Finally, Section VIII concludes.
II Related Work
@@ II-A The Concept of Fairness
Fairness is a complex, context-dependent, and ambiguously defined concept [13, 15, 12]. Relevant for our work are individual attitudes, biases, and legality, morality and ethics [19].
II-A1 Individual attitudes
@@
This aspect includes distributive, procedural, and interactional justice perceptions [20]. We notice that justice perceptions are often used interchangeably with fairness. This is partially in line with the literature, with some adding that fairness is downstream of justice [21, 22]. Others also split interactional justice into interpersonal and informational justice [22]. Because privacy policies are equally accessible to every user, we do not need to consider allocation aspects (distributive fairness), and focus on procedural and informational fairness.
@@
Procedural fairness [23] means procedures that are unbiased and unaffected by ideology. They should be representative of the individuals involved, and rely on accurate information. Fair processes should respectfully treat those affected by them. Assessing procedural fairness can be based on the extent to which procedures “suppress bias, create consistent allocations, rely on accurate information, are correctable, represent the concerns of all recipients, and are based on moral and ethical standards” [20].
@@
Informational fairness [22, 24] is about clearly, consistently and reasonably explained processes and whether the information provided is suited to individuals’ needs. Informational fairness addresses how complete the disclosed information is and how readable and comprehensible it is to its audience. Informational fairness is closely related to transparency, which can be seen as an aspect of procedural fairness [25]. While completeness is use-case-specific, readability [22] refers to linguistic approaches.
II-A2 Biases
@@
Embedded meanings in language can be biased [19]. A bias is a “dynamic and social and not [just] a statistical issue” [26]. Biases impact fairness, because they can be harmful to specific groups [27]. A prominent bias in text corpora is the presence of stereotypes, which is addressed by representational fairness.
@@
Representational fairness [28] can be a source of harm, because language establishes power relationships and represents social identities [27]. Consequently, stereotypical associations of a demographic group represented in a bias can lead to discrimination. Discrimination is the unjustified difference in the treatment of individuals based on their membership in different social / demographic groups [29], and can be a source of unfairness [16]. Representational unfairness materializes in word embeddings as toxicity, stereotyping, or other forms of misrepresentation of demographic groups, and has received much attention [30, 31, 9, 27] already.
II-A3 Legality, morality and ethics
Equal treatment is a fundamental human right, as declared in anti-discrimination principles in the Universal Declaration of Human Rights (UDHR) [32], the European Convention on Human Rights (ECHR) [33], and respective regulations in national law, e.g., the US [29].
Morality and ethics play an important role in fairness [19, 12, 13]. Ethics can be split into descriptive ethics and normative ethics [34]. Descriptive ethics addresses society’s perception or intuition of what is moral or not. Normative ethics can follow a consequentialist, deontological or virtue ethics approach in establishing generalizable norms that help in making moral judgments. The consequentialist approach tries to quantify and weigh all the positive impacts of an action against its negative impacts on everyone affected. The deontological approach judges an action based on how ethically desirable the maxim behind it would be if it were generally applied. The minimization of (potential) harm can be seen as a desirable principle and in line with consequentialist ethics. As not every bias must induce harm [27], bias is not necessarily morally questionable. Virtue ethics claims all actions that align with virtues to be moral. Fairness could also be seen as a virtue, following virtue ethics. Morality and ethics issues could be seen as a deviation from fundamental legal frameworks [32, 33, 1]. Prior research has also aimed at principles to guide machine ethics [35], e.g., to establish socially acceptable moral preferences for autonomous driving.
II-B Machine Learning and NLP Approaches
@@
II-B1 Readability Metrics
Related metrics exist on three levels. On the word level, foreign words, anglicisms, long words, abbreviations, complex words, etc., are problematic for non-native speakers [4], elderly [5], dyslexics [6] or autistic people [7], or socioeconomically less-privileged [8]. Lexical ambiguity of words can be assessed by comparing domain-specific meanings of words with their common meaning, e.g., by using BERT word embeddings with the Bradley-Terry statistical model [36].
On the document level, the Flesch Reading Ease (FRE) [37] is a commonly used metric for readability [38]. However, FRE does not capture readability holistically [39]. Modern NLP-based approaches like TAASSC 2.0 [40] provide further metrics like surface indices, metrics for syntactic features, as well as metrics for semantics. Comparing the consistency and logic of semantics between sentences allows coherence metrics, e.g., DiscoScore [41], which measure thematic focus frequency and sentence connectivity.
Finally, metrics on the structural level consider that a better-structured text improves the understanding [42]. Assessing structure separately is recommended [43].
II-B2 Bias Metrics
@@
The identification of biases typically needs a list of descriptor terms. Such descriptor terms are word lists in specific demographic axes that capture the different groups that could be misrepresented. For example, HolisticBias [44] includes 600 descriptor terms across 13 different demographic axes such as ability, age, body type, nationality, politics, etc. Descriptor terms can be generated automatically, e.g., for assessing gender bias on legal documents [10]. Assessing semantics has become the standard practice in literature for assessing representational fairness [30, 45, 9, 10, 46, 11]. To measure a bias, the text in question is transformed into word embeddings, i.e., a numerical, vectorized representation of the words used. Classical approaches measure cosine similarity between the descriptor terms, utilizing word embeddings to detect biases [45, 11, 30, 46]. A popular metric is WEAT [30], improved in its robustness to SAME [46].
@@
Other approaches are based on text features (e.g., using all-caps words or bold formatting) or semantic features (e.g., sarcasm or insinuation of dishonesty or deceitfulness). For example, see [47], a solution for the GermEval 2021 Shared Task on the Identification of Toxic Engaging, and Fact-Claiming Comments. This, however, can be biased in itself and needs bias mitigation [48].
II-C Morality and Ethics Classification
@@
Ethics might question the general morality of processes [49]. A judgment could, for instance, be derived following a set of principles / virtues like autonomy, nonmaleficence, beneficence, and justice [50]. However, research on classifying morality or ethics mainly focuses on moral stance prediction [51].
@@
Modern Large Language Models (LLM), be it open-source like Mixtral-8x7B [52], be it closed-source like GPT4turbo [53], show some cross-domain understanding [54], capabilities of reasoning [55] and some understanding of the concepts of morality and ethics [56]. To evaluate ethics and morality with them, modern prompt engineering approaches [57] are required.
II-D Evaluation of Privacy Policies
We are not aware of automated approaches to assessing the fairness of privacy policies. However, some machine learning and NLP approaches have already been applied. For example, research shows that privacy policies tend to be lengthy and use inaccessible language [58, 38]. For these reasons, approaches summarize key aspects of privacy policies with NLP [59, 60]. Another example is the assessment of the completeness and compliance of privacy policies regarding the GDPR [61, 62, 63] to mitigate a lack of transparency [2]. The most promising approach [61] first identifies informational requirements involving 56 metadata types relevant to the completeness of a policy. Based on that, it identifies 23 criteria to check for policies. The completeness assessment via Machine Learning (ML) and NLP is based on semantic similarity and word embeddings. To the best of our knowledge, representational fairness has not yet been assessed for privacy policies. However, representational fairness has been successfully assessed in loosely related fields of the legal domain [9, 10, 11] or more broadly on textual data [45, 30, 46].
@@ II-D Evaluation of Privacy Policies
Assessing morality and ethics requires looking into what policies enforce. Judging ethics is a field that is typically addressed by ethics councils. An ethics study of privacy policies has found that they “obfuscate, enhance and mitigate unethical data handling practices and use persuasive appeals to increase companies’ trustworthiness” [64]. This study is not in the context of the GDPR. However, the criteria of the GDPR can be used as a first step to evaluate the vulnerability and risks of a privacy policy. A prominent approach is to use a Support Vector Machine to sentence-wise quantify the degree of policy compliance with the data protection goals [63]. Most approaches do not make their data and code available.
III Problem Statement
@@
III-A Legal Foundation
Art. 12 of the Universal Declaration of Human Rights (UDHR) [32] and Art. 8 of the European Convention on Human Rights (ECHR)  [33] emphasize the right to privacy. The Privacy Guidelines [14], established in 1980, by the Organization for Economic Co-operation and Development (OECD), provide a framework for data privacy protection that 38 of the most developed countries have agreed upon, and use it as a reference for legislation. The guidelines require that “data should be obtained by lawful and fair means”, and implicitly reference various fairness concepts to establish a balanced approach to privacy protection.
The OECD Privacy Guidelines have been widely adopted. For example, the Asia-Pacific Economic Cooperative mirrors the Guidelines for its 21 member states [65]. In the USA and Canada, the OECD Guidelines are the basis of the Generally Accepted Privacy Principles (GAPP) [66], which were developed by the American Institute of CPAs and the Canadian Institute of Chartered Accountants for practical implementations. The GAPP are “based on internationally known fair information practices.” [66]
In the European Union, the GDPR [1] implements these guidelines into EU law. Chap. 2 GDPR specifies the openness and participation principles, which address informational fairness. Chap. 5 GDPR requires “ensuring that there is no unfair discrimination”. To this end, the GDPR requires any organization that manages personal data to declare in a privacy policy how it collects, shares, or makes use of its users’ data [2]. In particular, Art. 5(1) GDPR enforces fairness as a key principle for privacy policies [67]. Recital 39 highlights the close link to transparency and legality of this notion of fairness.
III-B Dimensions of Fairness
In the following, we focus on the GDPR. The GDPR and its legal foundation make three dimensions of fairness explicit:
Representational Fairness: Art. 1 UHDR states that equal treatment is a fundamental human right. Art. 2 UHDR and Art. 14 ECHR explicitly forbids discrimination on any ground such as sex, race, colour, language, religion, political or other opinion, national or social origin, association with a national minority, property, birth or other status.”

### v2

@@
Abstract.
Privacy policies are expected to inform data subjects about their data protection rights and should explain the data controller’s data management practices. Privacy policies only fulfill their purpose, if they are correctly interpreted, understood, and trusted by the data subject. This implies that a privacy policy is written in a fair way, e.g., it does not use polarizing terms, does not require a certain education, or does not assume a particular social background.
We outline our approach to assessing fairness in privacy policies. We identify from fundamental legal sources and fairness research, how the dimensions informational fairness, representational fairness and ethics / morality are related to privacy policies. We propose options to automatically assess policies in these fairness dimensions, based on text statistics, linguistic methods and artificial intelligence. We conduct initial experiments with German privacy policies to provide evidence that our approach is applicable. Our experiments indicate that there are issues in all three dimensions of fairness. This is important, as future privacy policies may be used in a corpus for legal artificial intelligence models.
@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 10th ACM International Workshop on Security and Privacy Analytics; June 21, 2024; Porto, Portugalbooktitle: Proceedings of the 10th ACM International Workshop on Security and Privacy Analytics (IWSPA ’24), June 21, 2024, Porto, Portugaldoi: 10.1145/3643651.3659890isbn: 979-8-4007-0556-4/24/06ccs: Security and privacy Social aspects of security and privacyccs: Social and professional topics Privacy policies
1. Introduction
The General Data Protection Regulation (GDPR) (European Union, 2016) requires any organization, that manages personal data, to publish a privacy policy. A policy should make transparent how personally identifiable information is collected, shared, or used (Zaeem and Barber, 2020). Privacy policies fulfill an important social task: They balance the information deficit between the data subject and holder, and create trust (Starke et al., 2022).
To ensure that privacy policies are properly understood, perceived, and accepted, they must be written in a fair way. This includes multiple dimensions. Informational Fairness is not just about what information is communicated, but also how it is communicated (Boudjella et al., 2017; Trzepla, 2019; Rello and Baeza-Yates, 2017; Evans et al., 2014; Aikens and Barbarin, 2008). A negative example would be a policy consisting of complex legal phrases, which discriminates against people with dyslexia and non-native speakers. Representational Fairness is about biases towards certain groups represented in texts (Rice et al., 2019; Baker Gillis, 2021; Gumusel et al., 2022). For example, a policy could discriminate against women by only using male word forms. Ethics and Morality are at the core of fairness (Hooker, 2005; Schwöbel and Remmers, 2022) and to some extent captured by legal frameworks (OECD, 2013; European Union, 2016). A negative example would be a policy, that tries to take exclusive rights to utilize the user’s personal data. However, many different definitions of fairness exist (Verma and Rubin, 2018; Mehrabi et al., 2021), some of them are contradictory (Chouldechova, 2017; Defrance and De Bie, 2023), and to the best of our knowledge, none of them is tailored for privacy policies. It is also unclear, how well methods from natural language processing or artificial intelligence allow evaluating privacy policies for such issues. Thus, our research question is as follows:
How can we automatically assess informational fairness, representational fairness, and ethics / morality of privacy policies?
In this paper, we relate prominent definitions and concepts of fairness and bias to privacy policies. We propose approaches for automatically evaluating privacy policies, and we test this with selected policies. We make three contributions:
@@ 1. Introduction
Paper structure: Section 2 reviews related work. In Section 3, we derive our problem statement from legal requirements. The Sections 4, 5 and 6 explain how three dimensions of fairness can be automatically assessed. Section 7 summarizes potential applications and use cases of our approach. Finally, Section 8 concludes.
2. Related Work
@@ 2.1. The Concept of Fairness
Fairness is a complex, context-dependent, and ambiguously defined concept (Schwöbel and Remmers, 2022; Verma and Rubin, 2018; Hooker, 2005). Relevant for our work are individual attitudes, biases, and legality, morality and ethics (Landers and Behrend, 2022).
2.1.1. Individual attitudes
@@
This includes distributive, procedural, and interactional justice perceptions (Greenberg, 1990). We notice that justice perceptions are often used interchangeably with fairness. Others also split interactional justice into interpersonal and informational justice (Colquitt and Rodell, 2015). Because privacy policies are equally accessible to every user, we do not need to consider allocation aspects (distributive fairness), and focus on procedural and informational fairness.
@@
Procedural fairness (Doherty and Wolak, 2012) means unbiased and non-ideological procedures. They should represent individuals involved, and rely on accurate information. Fair processes should respectfully treat affected individuals. Assessing procedural fairness can be based on how procedures “suppress bias, create consistent allocations, rely on accurate information, are correctable, represent the concerns of all recipients, and are based on moral and ethical standards” (Greenberg, 1990).
@@
Informational fairness (Colquitt and Rodell, 2015; Schoeffer et al., 2022) is about clearly, consistently and reasonably explained processes and whether the information provided is suited to individuals’ needs. Informational fairness addresses how complete and specific the disclosed information is and how readable and comprehensible it is to its audience. Informational fairness is closely related to transparency, which can be seen as an aspect of procedural fairness (Lee et al., 2019).
2.1.2. Biases
@@
Embedded meanings in language can be biased (Landers and Behrend, 2022). A bias is a “dynamic and social and not [just] a statistical issue” (Ntoutsi et al., 2020). Biases impact fairness, because they can be harmful to specific groups (Blodgett et al., 2020). A prominent bias in text corpora is the presence of stereotypes, which is addressed by representational fairness.
@@
Representational fairness (Abbasi et al., 2019) can be a source of harm, because language establishes power relationships and represents social identities (Blodgett et al., 2020). A bias with stereotypical associations of a demographic group can lead to discrimination. Such a demographic or social group is called protected group. Discrimination is the unjustified difference in the treatment of individuals based on their membership in protected (sub)groups (Edenberg and Wood, 2023), and can be a source of unfairness (Mehrabi et al., 2021; Wachter et al., 2021). Representational unfairness materializes in word embeddings as toxicity, stereotyping, or other forms of misrepresentation of protected groups, and has received much attention (Caliskan et al., 2017; Papakyriakopoulos et al., 2020; Rice et al., 2019; Blodgett et al., 2020; Garg et al., 2018; Gumusel et al., 2022; Schröder et al., 2021) already.
2.1.3. Morality and ethics
Morality and ethics play an important role in fairness (Landers and Behrend, 2022; Hooker, 2005; Schwöbel and Remmers, 2022). Privacy ethics investigates access provided to others to personal information and how much control one has regarding one’s information being collected, stored, and used by others (DeCew, 1986). In the context of modern technologies, it discusses complex privacy trade-offs and power relationships between the data holder and data subject (Acquisti et al., 2015; Clifford and Ausloos, 2018). Ethics and law share a coordinating function and complement each other (Rochel, 2021). It uses values and norms to interpret law. The latter addresses soft ethics which (Floridi, 2018) distinguishes from hard ethics which shape legislation like the GDPR and are embedded in it.
2.2. Machine Learning and NLP Approaches
@@
2.2.1. Readability Metrics
Related metrics exist on three levels. On the word level, foreign words, anglicisms, long words, abbreviations, complex words, etc., are problematic for non-native speakers (Boudjella et al., 2017), elderly (Trzepla, 2019), dyslexics (Rello and Baeza-Yates, 2017) or autistic people (Evans et al., 2014), or socioeconomically less-privileged (Aikens and Barbarin, 2008). Lexical ambiguity of words can be assessed by comparing domain-specific meanings of words with their common meaning, e.g., by using BERT word embeddings with the Bradley-Terry statistical model (Liu et al., 2022). On the document level, the Flesch Reading Ease (FRE) (Flesch, 1948) is a commonly used metric for readability (Becher and Benoliel, 2021). However, FRE does not capture readability holistically (Crossley et al., 2017). Modern NLP-based approaches like TAASSC 2.0 (Kyle et al., 2021) provide metrics on surface indices, syntactic features, and semantics. Comparing the consistency and logic of semantics between sentences allows coherence metrics, e.g., DiscoScore (Zhao et al., 2023), which measure thematic focus frequency and sentence connectivity. Finally, metrics on the structural level consider that a better-structured text improves the understanding (Cocklin et al., 1984). Assessing structure separately is recommended (Power et al., 2003).
2.2.2. Bias Metrics
@@
The identification of biases typically needs a list of descriptor terms. Such descriptor terms are word lists in specific demographic axes that capture the different groups that could be misrepresented. For example, HolisticBias (Smith et al., 2022) includes 600 descriptor terms across 13 different demographic axes such as ability, age, body type, nationality, politics, etc. Descriptor terms can be generated automatically, e.g., for assessing gender bias on legal documents (Baker Gillis, 2021). Assessing semantics has become the standard practice in literature for assessing representational fairness (Caliskan et al., 2017; Garg et al., 2018; Rice et al., 2019; Baker Gillis, 2021; Schröder et al., 2021; Gumusel et al., 2022). To measure a bias, the text in question is transformed into word embeddings, i.e., a numerical, vectorized representation of the words used. Classical approaches measure cosine similarity between the descriptor terms, utilizing word embeddings to detect biases (Garg et al., 2018; Gumusel et al., 2022; Caliskan et al., 2017; Schröder et al., 2021). A popular metric is WEAT (Caliskan et al., 2017), improved in its robustness to SAME (Schröder et al., 2021).
@@
Other approaches are based on text features (e.g., using all-caps words or bold formatting) or semantic features (e.g., sarcasm or insinuation of dishonesty or deceitfulness). For example, see (Risch et al., 2021), a solution for identifying toxic engaging, and fact-claiming comments. This, however, can be biased in itself and needs bias mitigation (Garg et al., 2023).
2.2.3. Morality and Ethics Classification
@@
Ethics might question the general morality of processes (Wallimann-Helmer et al., 2021). A judgment could be derived following a set of principles / virtues like autonomy, nonmaleficence, beneficence, and justice (Beauchamp and Childress, 2001). However, research on classifying morality or ethics focuses on moral stance prediction (Pavan et al., 2020).
@@
Modern Large Language Models (LLM), like Mixtral-8x7B (Jiang et al., 2024) or GPT4turbo (OpenAI, 2023), show some cross-domain understanding (Hendrycks et al., 2020), capabilities of reasoning (Bubeck et al., 2023) and some understanding of the concepts of morality and ethics (Pock et al., 2023). To evaluate ethics and morality with them, modern prompt engineering approaches (Nori et al., 2023) are required.
2.3. Evaluation of Privacy Policies
Fairness in privacy policies has been manually assessed in the mobile health domain (Benjumea et al., 2020). A rather narrow definition of fairness based on items of the 14 content requirements stated in Article 13 of the GDPR is embraced. We are unaware of an automated approach to assessing fairness of privacy policies independent of the application domain. Research shows that privacy policies tend to be lengthy and use inaccessible language (Bartelt and Buchmann, 2024; Becher and Benoliel, 2021). For these reasons, approaches summarize or tag key aspects of privacy policies with NLP and machine learning (Harkous et al., 2018; Nokhbeh Zaeem et al., 2020; Tesfay et al., 2018b). The completeness and compliance of privacy policies regarding the GDPR (Contissa et al., 2018; Amaral et al., 2021; Tesfay et al., 2018a; Sánchez et al., 2021; Vanezi et al., 2021; Xiang et al., 2023; Torre et al., 2020; Elluri et al., 2021) can be automatically assessed. Closest to our work is Claudette (Contissa et al., 2018), which assesses GDPR compliance based on completeness of provided information, substantive compliance of a policy, and clarity of expression. Assessing fairness of policies goes beyond assessing their compliance with law. Our assessment of informational fairness is also concerned with completeness of provided information, assessing readability in our approach also tries to identify vague language but, further, addresses issues like readability for protected groups. Our approach assesses representational fairness, which Claudette does not cover, and ethics, which Claudette partially covers with substantive compliance. The most promising approach for completeness of a policy (Amaral et al., 2021) first identifies informational requirements involving 56 metadata types relevant to the completeness of a policy. Based on that, it identifies 23 criteria to check for policies. The completeness assessment via Machine Learning (ML) and NLP is based on semantic similarity and word embeddings. Other approaches provide criteria or templates for more user-friendly and understandable policies (Renaud and Shepherd, 2018; Feng et al., 2021).
To the best of our knowledge, representational fairness has not yet been assessed for privacy policies. However, representational fairness has been successfully assessed in loosely related fields of the legal domain (Rice et al., 2019; Baker Gillis, 2021; Gumusel et al., 2022) or more broadly on textual data (Garg et al., 2018; Caliskan et al., 2017; Schröder et al., 2021).
@@ 2.3. Evaluation of Privacy Policies
Assessing morality and ethics requires looking into what policies enforce. Judging ethics is a field that is typically addressed by ethics councils. Privacy policies have been found to obfuscate unethical data handling practices and use persuasive language (Pollach, 2005). However, the criteria of the GDPR can be used as a first step to evaluate the risks of a privacy policy. For example, policies have been found to lack specificity regarding data use practices (Zaeem and Barber, 2020) and lack protection of minors as well as a clear communication of changes made to their policy (Contissa et al., 2018). A prominent approach to identifying risk is to use a Support Vector Machine to sentence- or paragraph-wise quantify the degree of policy compliance with the data protection goals (Contissa et al., 2018; Sánchez et al., 2021). Most approaches do not make their data and code available.
3. Problem Statement
@@
3.1. Legal Foundation
Equal treatment is a fundamental human right, as declared in anti-discrimination principles in the Universal Declaration of Human Rights (UDHR) (UN, 1948), the European Convention on Human Rights (ECHR) (Council of Europe, 1950), and respective regulations in national law, e.g., the US (Edenberg and Wood, 2023). Art. 12 of the UDHR and Art. 8 of the ECHR emphasize the right to privacy. The Privacy Guidelines (OECD, 2013) of the Organization for Economic Co-operation and Development (OECD) provide a framework for data privacy protection that is widely used as a reference for legislation. The guidelines require that “data should be obtained by lawful and fair means”, and implicitly reference various fairness concepts to establish a balanced approach to privacy protection. The OECD Privacy Guidelines have been widely adopted. For example, the 21 member states of the Asia-Pacific Economic Cooperative mirror the Guidelines (APEC, 2015). In the USA and Canada, the OECD Guidelines are the basis of the Generally Accepted Privacy Principles (GAPP) (AICPA and CICA, 2006). The GAPP are “based on internationally known fair information practices” (AICPA and CICA, 2006). In the European Union, the GDPR (European Union, 2016) implements these guidelines into EU law. Chap. 2 GDPR specifies the openness and participation principles, which address informational fairness. Chap. 5 GDPR requires “ensuring that there is no unfair discrimination”. Art. 5(1) GDPR enforces fairness as a key principle for privacy policies (Clifford and Ausloos, 2018; Malgieri, 2020).
3.2. Dimensions of Fairness
Legal foundations motivate our three dimensions of fairness:
Representational Fairness: Art. 1 UHDR states that equal treatment is a fundamental human right. Art. 2 UHDR and Art. 14 ECHR explicitly forbids discrimination on any ground”.